OpenStudy (gabby1214):

Help with solving another equation 8x=15 the solution is x=? ^^(an integer or a simplified fraction)

OpenStudy (chantysquirrel1129**):

Well, depends on what the question is asking for c:

OpenStudy (chantysquirrel1129**):

Do you know what the first step is? :3

OpenStudy (anonymous):

X=15/8

OpenStudy (anonymous):

Simplified is 1 7/8

OpenStudy (anonymous):

X=15/8

OpenStudy (anonymous):

x=15/8.